OpenStudy (anonymous):

The perimeter of a square is the product of four and the length of a side. Which of the following function rules could be used to find the perimeter of a square, represented by y, given the length of a side, represented by x?

OpenStudy (chrisdbest):

So what are the answer choices??

OpenStudy (anonymous):

My bad A.x = 4y B.y = 4x C.y = x + 4 D.x = y + 4

OpenStudy (chrisdbest):

Thats easy, Its B

OpenStudy (anonymous):

Alright thanks

OpenStudy (chrisdbest):

"The product of 4 and the length of a side" meaning 4 and a number/variable. The perimeter of a square is represented by y, so Y=4x

OpenStudy (chrisdbest):

Its ok

OpenStudy (chrisdbest):

plz medal